OpenStudy (anonymous):

What is the y-intercept of the line with the equation? y=4x+2

OpenStudy (kiakat):

2

OpenStudy (anonymous):

Ok @riverrrr do you know how to find the y-intercept of an equation?

OpenStudy (anonymous):

To find a y-intercept, set your x = 0 and solve for y. y = 4x + 2 y = 4(0) + 2 y = 2